If f(x) =10x-A/x-9 and (f) (-5) =0 , then what is value of A and where is f not defined?
Answer by rothauserc(4718)

f(x) = 10x-A/x-9 and (f) (-5) = 0
0 = (10*(-5) - A) / (-5 -9)
0 = -50 -A / -14
multiply both sides of = by -14
0 = -50 -A
-A = 50
A = -50
check answer
0 = (10*(-5) - (-50)) / (-5 -9)
0 = 0 / 14
0 = 0
note that the function f(x) is not defined at x = 9 since division by 0 is not defined